OpenStudy (anonymous):

if 7^(-x)=2 then what is 7^(3x)?

hartnn (hartnn):

first lets find the value of 7^x a^{-b} = 1/a^b so, 7^{-x} = 1/7^x =2 so, 7^x =...?

OpenStudy (anonymous):

2/7?

hartnn (hartnn):

where did 7 come from ? 1/7^x = 2 so, 7^x = 1/2 got this ?

OpenStudy (anonymous):

ohhok ! yes got it

OpenStudy (anonymous):

so then how would we get 7 to 3x?

hartnn (hartnn):

now 7^(3x) = \((7^x)^3=.... ?\)

OpenStudy (anonymous):

would it be 1/2 to 3?

OpenStudy (anonymous):

1/8?

hartnn (hartnn):

thats absolutely correct! good :)

OpenStudy (anonymous):

thank you so much:)

hartnn (hartnn):

welcome so much:)